# Alpine: An adaptive language-agnostic pruning method for language models for code

### v1

@@ Abstract
Language models of code have demonstrated state-of-the-art performance across various software engineering and source code analysis tasks. However, their demanding computational resource requirements and consequential environmental footprint remain as significant challenges. This work introduces Alpine, an adaptive programming language-agnostic pruning technique designed to substantially reduce the computational overhead of these models. The proposed method offers a pluggable layer that can be integrated with all Transformer-based models. With Alpine, input sequences undergo adaptive compression throughout the pipeline, reaching a size that is up to less their initial size, resulting in significantly reduced computational load. Our experiments on two software engineering tasks, defect prediction and code clone detection across three language models CodeBert, GraphCodeBert and UniXCoder show that Alpine achieves up to a 50% reduction in FLOPs, a 58.1% decrease in memory footprint, and a 28.1% improvement in throughput on average. This led to a reduction in CO2 by up to %. Importantly, it achieves the reduction in computation resources while maintaining up to 98.1% of the original predictive performance. These findings highlight the potential of Alpine in making language models of code more resource-efficient and accessible while preserving their performance, contributing to the overall sustainability of adopting language models in software development. Also, it sheds light on redundant and noisy information in source code analysis corpora, as shown by the substantial sequence compression achieved by Alpine.
I Introduction
@@
Transformer-based language models [1] were initially conceived for natural language understanding. Under the naturalness hypothesis [2], they were applied to solve diverse software engineering (se) tasks with remarkable accuracy. These tasks include code generation [3, 4, 5, 6], quality assurance [7, 8, 9, 10], maintenance [11, 12, 13, 14], and requirements engineering [15, 16].
@@
Though the language models have shown exceptional promise, fine-tuning or running inference on these models require significant computational resources. Moreover, the demand of the resources is continuously increasing with the size of the language models. This entails that their training grows correspondingly, leading to increased usage of cloud computing services which operate on a pay-per-use basis. This inflates the direct financial investment for researchers and average users with consumer-grade hardware. In addition to the financial impact, the carbon footprint of training and maintaining these models is a growing environmental concern as they contribute to higher CO2 emissions [17, 18, 19]. These issues underscore the urgency for more resource efficient algorithms.
@@
Shi et al. [20] have proposed Compressor as a step towards this direction. Compressor combines task-specific knowledge distillation [21] and evolutionary search of network architecture to generate a configuration that yields a smaller model. Task-specific knowledge distillation often falls short in terms of task-specificity and computational overhead. The distilled student models are typically optimized for a specific task or domain, limiting their generalization abilities. Moreover, the process of task-specific knowledge distillation itself can be computationally expensive, as it involves fine-tuning the teacher model either way, generating predictions or soft labels, pre-training, and fine-tuning the student model.
@@
Similarly, DietCode [22] attempts to achieve better resource utilization by assigning importance scores to tokens and statements based on attention weights, and then selecting the most important statements to form a new, condensed input. However, it has two main limitations. First, it relies on the peculiarities of task’s programming language, such as specific syntax, semantics, and constructs. This reliance makes extending the method to all programming languages difficult without overhead and customization. Second, it assumes that the attention weights remain static and applicable across different corpora and tasks. However, this is not the case given that these attention weights are derived from trainable parameters, which are updated when the language model is fine-tuned on other tasks.
@@
To address the issues above, we introduce Alpine, a programming-language-independent pruning method that can be directly integrated into any Transformer-based model. As sequences go through the layers, Alpine prunes unimportant tokens that lay within a configurable range, reducing input size, and offering a considerable speedup during fine-tuning and inference. We evaluate Alpine on two software engineering tasks defect prediction and code clone detection using three language models, CodeBert, GraphCodeBert and UniXCoder. Our approach allows these models to use less computation, measured by the number of floating point operations (flops) and memory footprint, with minimal loss in performance and even slight improvement in some scenarios.
@@
We conduct an analysis to quantify the computational requirements, measured by the flops, of the individual components within the Transformer model. The analysis helps us identify the bottleneck components and understand the changes in computational complexity with respect to the input sequence length.
@@
Replication package: Our replication package including source code and data is available online [23].
II Background and Motivation
II-A Transformer Architecture
The Transformer architecture, introduced by Vaswani et al. [1] for natural language processing (nlp), is a sequence-to-sequence model composed of two components: an encoder and a decoder. The focus of this work is on the encoder-only variant, one of the most used variants in SE-related tasks [24]. For the rest of the paper, we refer to a Transformer encoder layer by a Transformer layer for convenience.
Input Representation: The input fed into the Transformer layer consists of the tokenized sequence and an attention mask . includes a special [CLS] token, whose representation is used for classification or ranking tasks, and a [SEP] to indicate the end of the sequence, followed by padding tokens [PAD], if necessary111Appending two sequences is also another format that can be used as input: I.
@@ II-A Transformer Architecture
Main blocks: The main blocks of this layer are the Multi-head Self-Attention (mha) which is composed of attention heads, and the Position-wise Feed-forward network (ffnn):
Where is the scaled dot attention function introduced by Vaswani et al. [1], and is the input sequence. The derivation of Equation 1 is taken from the work of Kobayashi et al. [25] and Saad and Sharma [26]. Next, the encoder layer applies the ffnn to each position of the sequence independently. It is a subnetwork composed of two linear layers with the application of ReLU in between.
Note that, as shown above in the two equations, the Transformer layer employs residual connections followed by a layer normalization after each block to facilitate the flow of information and stabilize the training process.
II-B Pruning
II-C Motivation
II-C1 Impact on Energy Consumption
Language models for code require a significant amount of computational power to train and run. The number of parameters in these models has increased from millions to billions [29, 30], necessitating the use of high-performance computing systems and specialized hardware, such as graphics processing units (gpus) and tensor processing units (tpus). The training of such models involves processing large amounts of data, often in the order of terabytes. This requires not only powerful processors but also high-bandwidth memory and storage systems. As a result, the energy consumed during the training phase is substantial, as it involves running the models for extended periods that can reach the magnitude of days. Table LABEL:tab:model_training_time illustrates the training time of different models of various sizes.
@@ II-C1 Impact on Energy Consumption
| UniXCoder [31] | 125M | 8 days |
| CodeT5base [32] | 220M | 12 days |
| InCoder [33] | 6.7B | 24 days |
The energy consumption of language models is directly proportional to their computational requirements. As the models become more complex and the amount of data they process increases, the energy needed to power the computing systems also rises. This results in a higher carbon footprint of these models. Wang et al. [19] have shown that the carbon footprint of pre-training the bert model [34] on a machine with one NVIDIA Quadro RTX gpu is kg CO2. This is equivalent to the same amount emitted by driving a vehicle for a distance of kilometers222Wang et al. [19] report the distance in miles [19]. Moreover, we are witnessing a notable trend of models increasing in size leading to even higher environmental impact. For instance, recently, Lozhkov et al. [35] released StarCoder 2, a family of language models for code with B to B parameters. They reported that the smallest model has resulted in kg of CO2 emission during training, excluding the fine-tuning stage.
II-C2 Computational Cost of the Transformer Layer
The pointwise feed-forward network represents a significant memory bottleneck, especially in comparison to the multi-head attention layer. This section explores the computational complexity and memory requirements of both the mha and ffnn layers, highlighting how the latter’s larger dimension contributes to such a bottleneck.
The computational cost of a neural layer can be measured in terms of the number of floating-point operations (flops) required for each layer. We calculate the theoretical operations count, i.e., the upper bound where we have a sequence with the maximum number of tokens. The flops counting rules introduced by Hunger [36] for matrix multiplications are used in our work. Specifically, the rule stating that the number of flops of a matrix multiplied with another matrix is .
@@ II-C2 Computational Cost of the Transformer Layer
The next operation in the mha layer involves the multiplication of the attention probability matrices of each head obtained in Equation 7 with the value matrix that was calculated from one of the linear projections performed earlier.
Finally, the outputs of all attention heads are concatenated and projected through a linear layer. The concatenation operation requires 0 operations, hence, the only flops performed here stem the matrix multiplication of the concatenated matrix with .
Summing the Equations 3-9, we get the following formula to determine the flops count of the attention layer.
FLOPs of the ffnn: The ffnn takes the output of the mha’s layer as input, scales it to a higher dimension than the mha’s, applies a gelu transformation [37], and finally scales back to the original hidden dimension333While the Transformer layer was stated earlier to use ReLU activation, the models used in this work replace it with the gelu function..
@@
The ffnn’s dimension is usually set to be higher than the mha’s. For instance, in the models that we have chosen for our experiments, . Using this fact, and the Equations 11 and 12 the total number of flops is,
Although Equations 10 and 13 show the computational cost of mha increases quadratically, it uses fewer flops when the sequence length is small compared to the ffnn. Figure 1 illustrates the growth of the number of flops of both layers when sequences’ lengths range between and using the above formulae444We also provide an analytical proof in the replication package.. Such range is the maximal range for many language models used to solve SE tasks. This is further demonstrated empirically when measuring the number of flops performed by both of these layers in CodeBert [38] using the test set of the Devign [39] dataset. Thus, a potential way to reduce such complexity is to reduce the length of the sequence before feeding it to the ffnn layer.
III Approach
@@
Next, we present the pruning procedure adopted in this study in each Transformer layer and elaborate on the design choices.
@@
As shown in Figure 2 and Algorithm 2, the pruning layer is inserted between the mha and ffnn layers. As highlighted in Figure 1, the feed-forward layer represents a bottleneck in terms of computational costs and memory consumption [40] caused by the projection of the attention layer’s output to a higher dimension in the ffnn. Hence, performing pruning before this layer would reduce the effect of such a bottleneck.
@@
Algorithm 1 presents the steps to perform the pruning. The first step is to assign a score to each token to determine its importance within the sequence. In this work, we use attention based scoring [28], which quantifies this importance using the attention probabilities output by the mha layer. Specifically, in L4-11 we take the mean across the attention matrices of the attention heads of the mha layer and then perform a column-wise mean to get a distribution of importance scores as illustrated by Equation 14, where is the number of attention heads, is the attention matrix of an attention head and is the list of tokens excluding the special tokens [CLS], [SEP], and [PAD].
@@
The reason we exclude the [CLS] and [SEP] tokens from pruning is two-fold. First, the [CLS] token representation is extracted from the output of the final transformer encoder layer and fed into a linear classification layer as done previously [20, 41] Hence, such a token should be kept throughout the encoder pipeline. Second, it has been shown that these two tokens usually receive the highest attention scores [42, 25, 26] in natural and programming languages, which would skew the importance scores distribution. As for the padding token, its attention score is 0, which again would affect the score distribution, especially for shorter sequences.
@@
Once we obtain the importance scores distribution, we keep the tokens that are within a specific range. We set the range to be , where and are the mean and standard deviation of the importance scores distribution, and is a hyperparameter that defines the width of the lower and upper bounds as shown in L12-16. The smaller gets, the tighter the window, the less tokens are kept. This hyperparameter controls the aggressiveness of pruning.
@@
Given the importance scores and the range , we create a new mask that indicates the tokens that should be pruned and those that should be kept (L17-20),
@@
Algorithm 2 shows the modified implementation of the Transformer layer. The highlighted lines L7 and L8 represent the modification that we have introduced. This also demonstrates the ease of integrating Alpine with minimal effort.
@@
Using the updated mask from the previous step, we remove the rows from the mha output matrix that correspond to the tokens to be pruned as shown in L8. In an alternative experiment setting, we merge the rows identified for pruning into one row rather than completely removing them. Our reason is to minimize the information loss incurred by pruning. Next, the matrix’s dimensions are reduced. We also calculate the final mask since this operation is performed on a batch of sequences to guarantee that all sequences are of equal length.
@@
We also experiment with three variants of pruning. The first one involves performing pruning across all layers of the encoder. In the second setting, it is performed only for even-indexed layers (i.e., ). Whereas the final one involves pruning at odd-indexed layers (i.e., ).
IV Experiments

### v2

# An Adaptive Language-Agnostic Pruning Method for Greener Language Models for Code

@@ Abstract.
Language models of code have demonstrated remarkable performance across various software engineering and source code analysis tasks. However, their demanding computational resource requirements and consequential environmental footprint remain as significant challenges. This work introduces Alpine, an adaptive programming language-agnostic pruning technique designed to substantially reduce the computational overhead of these models. The proposed method offers a pluggable layer that can be integrated with all Transformer-based models. With Alpine, input sequences undergo adaptive compression throughout the pipeline, reaching a size that is up to less their initial size, resulting in significantly reduced computational load. Our experiments on two software engineering tasks, defect prediction and code clone detection across three language models CodeBert, GraphCodeBert and UniXCoder show that Alpine achieves up to a 50% reduction in FLOPs, a 58.1% decrease in memory footprint, and a 28.1% improvement in throughput on average. This led to a reduction in CO2 emissions by up to %. Importantly, it achieves a reduction in computation resources while maintaining up to 98.1% of the original predictive performance. These findings highlight the potential of Alpine in making language models of code more resource-efficient and accessible while preserving their performance, contributing to the overall sustainability of their adoption in software development. Also, it sheds light on redundant and noisy information in source code analysis corpora, as shown by the substantial sequence compression achieved by Alpine.
1. Introduction
@@
Transformer-based language models (Vaswani et al., 2017) were initially conceived for natural language understanding. Under the naturalness hypothesis (Hindle et al., 2012), they were applied to solve diverse software engineering (se) tasks with remarkable accuracy. These tasks include code generation (Lai et al., 2023; Zeng et al., 2022a; Dibia et al., [n. d.]; Chen et al., 2022), quality assurance (Alqarni and Azim, 2022; Ciborowska and Damevski, 2023; Fatima et al., 2023; Tang et al., 2023), maintenance (Tian et al., 2023; Zhang et al., 2023; Le-Cong et al., 2023; Paul et al., 2023), and requirements engineering (Moharil and Sharma, 2023; Ezzini et al., [n. d.]).
@@
Though the language models have shown exceptional promise, fine-tuning or running inference on these models require significant computational resources. Moreover, the demand of the resources is continuously increasing with the size of the language models. This entails that their training grows correspondingly, leading to increased usage of cloud computing services which operate on a pay-per-use basis. This inflates the direct financial investment for researchers and average users with consumer-grade hardware. In addition to the financial impact, the carbon footprint of training and maintaining these models is a growing environmental concern as they contribute to higher CO2 emissions (Luccioni et al., 2023; Lacoste et al., 2019; Wang et al., 2023b). These issues underscore the urgency for more resource efficient algorithms.
@@
Shi et al. (Shi et al., 2023) have proposed Compressor as a step towards this direction. Compressor combines task-specific knowledge distillation (Hinton et al., 2015) and evolutionary search of network architecture to generate a configuration that yields a smaller model. Task-specific knowledge distillation often falls short in terms of task-specificity and computational overhead. The distilled student models are typically optimized for a specific task or domain, limiting their generalization abilities. Moreover, the process of task-specific knowledge distillation itself can be computationally expensive, as it involves fine-tuning the teacher model either way, generating predictions or soft labels, pre-training, and fine-tuning the student model.
@@
Similarly, DietCode (Zhang et al., 2022) attempts to achieve better resource utilization by assigning importance scores to tokens and statements based on attention weights and then selecting the most important statements to form a new, condensed input. However, it has two main limitations. First, it relies on the peculiarities of the task’s programming language, such as specific syntax, semantics, and constructs. This reliance makes extending the method to all programming languages difficult without overhead and customization. Second, it assumes that the attention weights remain static and applicable across different corpora and tasks. However, this is not the case given that these attention weights are derived from trainable parameters, which are updated when the language model is fine-tuned on other tasks.
LTP (Kim et al., 2022) covers some of DietCode’s limitations. It proposes a method that learns a threshold at each Transformer layer and tokens are pruned if their importance score is below such threshold. Concretely, the pruning is done by zeroing out the hidden representations of these tokens. However, such an approach does not effectively reduce computation. While the method assigns zero values to pruned tokens by multiplying the hidden representations with a mask111LTP’s implementation: https://t.ly/qpu4J, the sequence length remains unchanged. Consequently, these zero-valued representations still participate in subsequent matrix multiplications and other operations. In most hardware and software implementations, multiplications involving zero are not automatically optimized out. As a result, the same number of floating-point operations (FLOPs) are performed regardless of how many tokens are “pruned” leading to no actual reduction in computational cost. Finally, LTP requires multiple fine-tuning steps: in the first pass, it learns the thresholds as well as full model parameters, then fine-tunes the model’s weights in the second pass using those thresholds.
To tackle the aforementioned limitations, we present Alpine, a language-agnostic token pruning technique seamlessly integrated into any Transformer-based model. This approach involves placing a pruning module within the Transformer block. Alpine processes a code snippet, generating a set of importance scores for each token. Based on these scores and a configurable importance range, it completely eliminates tokens that exist outside the range. Consequently, the reduced input size yields considerable speed up in both fine-tuning and inference phases. Notably, Alpine operates without requiring supplementary code information, such as statement types, ensuring its programming-language independence. Furthermore, the importance scores are computed during the fine-tuning process, enabling it to adapt dynamically to various downstream Software Engineering (SE) tasks.
@@
We evaluate our proposed method on two software engineering tasks defect prediction and code clone detection using three language models, CodeBert, GraphCodeBert and UniXCoder. Our approach allows these models to use less computation, measured by the number of floating point operations (flops) and memory footprint, with minimal loss in performance and even slight improvement in some scenarios.
@@
We conduct an analysis to quantify the computational requirements of the individual components within the Transformer model measured by flops. The analysis helps us identify the bottleneck components and understand the changes in computational complexity with respect to the input sequence length.
@@
Replication package: Our replication package including source code and data is available online (Saad et al., [n. d.]).
2. Background and Motivation
In this section, we provide a brief primer on the Transformer architecture (Vaswani et al., 2017) and the implications of its computational requirements. Then, we investigate the number of operations performed by the main blocks of this architecture to identify a potential bottleneck within a certain class of Transformer-based models.
2.1. Transformer Architecture
The Transformer architecture, introduced by Vaswani et al. (Vaswani et al., 2017) for natural language processing (nlp), is a sequence-to-sequence model composed of two main components: an encoder and a decoder. This work focuses on the encoder-only variant, one of the most used variants in SE-related tasks (Hou et al., 2023b). We further elaborate in Section 6.3 how Alpine can be extended to encoder-decoder and decoder-only models. For the rest of the paper, we refer to a Transformer encoder layer by a Transformer layer for convenience.
Input Representation: The input fed into the Transformer layer consists of the tokenized sequence of tokens and an attention mask . includes a special [CLS] token, whose representation is used for classification or ranking tasks, and a [SEP] to indicate the end of the sequence, followed by padding tokens [PAD], if necessary222Appending two sequences is also another format that can be used as input: I.
@@ 2.1. Transformer Architecture
Main blocks: The main blocks of this layer are the Multi-head Self-Attention (mha) which is composed of attention heads, and the Position-wise Feed-forward network (ffnn). mha is defined as follows
where is the scaled dot attention function introduced by Vaswani et al. (Vaswani et al., 2017), and is the input sequence. The derivation of Equation 1 is taken from the work of Kobayashi et al. (Kobayashi et al., 2020) and Saad and Sharma (Saad and Sharma, 2023). After the mha, the ffnn block is applied to each position of the sequence independently. It is a subnetwork composed of two linear layers with the application of a non-linear activation function ReLU in between.
Note that, at the end of each block, the Transformer layer employs residual connections followed by a layer normalization to facilitate the flow of information and stabilize the training process.
2.2. Motivation
2.2.1. Impact on Energy Consumption
Language models for code require a significant amount of computational power to train and run. The number of parameters in these models has increased from millions to billions (Hou et al., 2023a; Zhao et al., 2023), necessitating the use of high-performance computing systems and specialized hardware, such as graphics processing units (gpus) and tensor processing units (tpus). The training of such models involves processing large amounts of data, often in the order of terabytes. This requires not only powerful processors but also high-bandwidth memory and storage systems. As a result, the energy consumed during the training phase is substantial, as it involves running these models for extended periods that can reach the magnitude of days. Table LABEL:tab:model_training_time illustrates the training time of different models of various sizes.
@@ 2.2.1. Impact on Energy Consumption
| UniXCoder (Guo et al., 2022) | 125M | 8 days |
| CodeT5base (Wang et al., 2021) | 220M | 12 days |
| InCoder (Fried et al., 2023) | 6.7B | 24 days |
The energy consumption of language models is directly proportional to their computational requirements. As the models become more complex and the amount of data they process increases, the energy needed to power the computing systems also rises. This results in a higher carbon footprint of these models. Wang et al. (Wang et al., 2023b) have shown that the carbon footprint of pre-training the bert model (Devlin et al., 2018) on a machine with one NVIDIA Quadro RTX gpu is kg CO2. This is equivalent to the same amount emitted by driving a vehicle for a distance of kilometers333Wang et al. (Wang et al., 2023b) report the distance in miles (Wang et al., 2023b). Moreover, we are witnessing a notable trend of models increasing in size leading to even higher environmental impact. For instance, recently, Lozhkov et al. (Lozhkov et al., 2024) released StarCoder 2, a family of language models for code with sizes ranging from B to B parameters. They reported that the smallest model has resulted in kg of CO2 emission during pre-training.
2.2.2. Computational Cost of the Transformer Layer
This section explores the computational complexity and memory requirements of both the mha and ffnn layers.
The computational cost of a neural layer can be measured in terms of the number of floating-point operations (flops) performed by such layer. We calculate the theoretical operations count, i.e., the upper bound when a sequence has the length of the maximum number of tokens. The flops counting rules introduced by Hunger (Hunger, 2005) for matrix multiplications are used in our work. Specifically, the rule stating that the number of flops of a matrix multiplied with another matrix is . Note that here and represent the dimensions of the matrices and , respectively.
@@ 2.2.2. Computational Cost of the Transformer Layer
The next operation in the mha layer is the multiplication of the attention probability matrices of each head obtained in Equation 8 with the value matrix that was calculated from one of the linear projections performed earlier.
Finally, the outputs of all attention heads are concatenated and projected through a linear layer. The concatenation operation requires 0 operations, hence, the only flops performed are from the matrix multiplication of the concatenated matrix with .
Summing the Equations 4-10, we get the following formula to determine the flops count of the attention layer.
FLOPs of the ffnn: The ffnn takes the output of the mha’s layer as input, scales it to a higher dimension than the mha’s, applies a gelu transformation (Hendrycks and Gimpel, 2016)444While the Transformer layer was stated earlier to use ReLU activation, the models used in this work replace it with the gelu function. Note that this does not affect the calculations performed above., and scales it back to the original hidden dimension.
@@
The ffnn’s dimension is usually set to be higher than the mha’s. For instance, in the models that we have chosen for our experiments, . Using this fact, and the Equations 12 and 13 the total number of flops is,
Although Equations 11 and 14 show that the computational cost of mha increases quadratically compared to the ffnn, it uses fewer flops when the maximum sequence length is . Such length is an upper bound for many language models such as CodeBert (Feng et al., 2020), GraphCodeBert (Guo et al., 2021), and UniXCoder (Guo et al., 2022) that are used to solve SE tasks. We now prove this statement analytically,
Lemma 0.
For every conceivable input sequence fed into a Transformer model with hyperparameters same as CodeBert, GraphCodeBert, and UniXCoder, it holds true that
Proof.
Let us fix , and consider the following function:
This function
is quadratic
has a maximum .
has two roots and
therefore or when . In other words, if the sequence length , then
For UniXCoder, GraphCodeBert, and CodeBert, we have that and . Thus,
Furthermore, the maximum sequence lengths are 1024, 512, and 512, for UniXCoder, GraphCodeBert, and CodeBert respectively. Therefore, it will consistently hold true that , leading to the conclusion that .
This is further demonstrated empirically when measuring the number of flops performed by both of these layers in CodeBert (Feng et al., 2020) using the test set of the Devign (Zhou et al., 2019) dataset as shown in Figure 1.
Therefore, reducing the sequence length before feeding it to the ffnn layer is a potential way to reduce such complexity.
3. Approach
@@
In this section, we present the pruning procedure adopted in this study in each Transformer layer and elaborate on the design choices.
@@
As shown in Figure 2 and Algorithm 2, the pruning layer is inserted between the mha and ffnn layers. As highlighted in Lemma 2.1, the feed-forward layer represents a bottleneck in terms of computational costs and memory consumption (Ganesh et al., 2021) caused by the projection of the attention layer’s output to a higher dimension in the ffnn. Hence, performing pruning before this layer would reduce the effect of such a bottleneck.
@@
Algorithm 1 presents the steps to perform the pruning. The first step is to assign a score to each token to determine its importance within the sequence. In this work, we use attention based scoring (Kim et al., 2022), which quantifies this importance using the attention probabilities output by the mha layer. Specifically, in Lines 4–11 we take the mean across the attention matrices of the attention heads of the mha layer and then perform a column-wise mean to get a distribution of importance scores as illustrated by Equation 15, where is the number of attention heads, is the attention matrix of an attention head and is the list of tokens excluding the special tokens [CLS], [SEP], and [PAD].
@@
The reason we exclude the [CLS] and [SEP] tokens from pruning is two-fold. First, the [CLS] token representation is extracted from the output of the final transformer encoder layer and fed into a linear classification layer as done previously (Shi et al., 2023; Lu et al., 2021). Hence, such a token should be kept throughout the encoder pipeline. Second, it has been shown that these two tokens usually receive the highest attention scores (Clark et al., 2019; Kobayashi et al., 2020; Saad and Sharma, 2023) in natural and programming languages, which would skew the importance scores distribution. As for the padding token, its attention score is 0, which again would affect the score distribution, especially for shorter sequences.
@@
Once we obtain the importance scores distribution, we keep the tokens that are within a specific range. We set the range to be , where and are the mean and standard deviation of the importance scores distribution, and is a hyperparameter that defines the width of the lower and upper bounds as shown in Lines 12-16. The smaller gets, the tighter the window, the less tokens are kept. This hyperparameter controls the aggressiveness of pruning.
Based on a preliminary investigation, we observed that the distribution of importance tokens follows roughly a leptokurtic distribution (i.e., distribution with low standard deviation and a high peak) at each layer. By removing tokens outside the defined range in such distribution, we can preserve the most informative and representative data while reducing computational complexity.
@@
Given the importance scores and the range , we create a new mask that indicates the tokens that should be pruned and those that should be kept (Lines 17-20),
@@
Algorithm 2 shows the modified implementation of the Transformer layer. The highlighted Lines 7 and 8 represent the modifications that we have introduced. This also demonstrates the ease of integrating Alpine with minimal effort.
@@
Using the updated mask from the previous step, we either remove the rows from the mha output matrix that correspond to the tokens to be pruned as shown in Line 8, or merge them into one row. By default, we do the merging to minimize the information loss incurred by pruning, especially that of tokens that are above the upper bound of the importance interval. Next, the matrix’s dimensions are reduced. We also calculate the final mask since this operation is performed on a batch of sequences to guarantee that all sequences are of equal length.
@@
We experiment with three variants of pruning. The first one involves performing pruning across all layers of the encoder. In the second setting, it is performed only for even-indexed layers (i.e., ). Whereas the final one involves pruning at odd-indexed layers (i.e., ).
4. Experiments